# WaveletGaussian: Wavelet-domain Diffusion for Sparse-view 3D Gaussian Object Reconstruction

### v2

@@ 5.2 Quantitative Results
We present the quantitative results in Table 1. Generally, compared to the closest baseline, GaussianObject [GaussianObject], our proposed method achieves a 0.3-0.5 dB increase in PSNR and cuts the overall training time roughly by 40%.
@@ 5.3 Ablation Studies
Table 2 presents ablation results. Firstly, we replace the LOO strategy, utilized by the baseline, with two variations of the random masking strategy. Different from the Online RM strategy presented in Section 4.2, the Offline RM strategy does not incorporate drifting masks. The former achieves better PSNR because the more diverse corruption patterns make $\mathcal{D}$ more robust. Both strategy outperform LOO in training time due to training a single $\mathcal{G}_{d}$, and without performance reductions. Having incorporated ORM, we then use wavelet diffusion (“wavelet-$\mathcal{D}$”) for novel view repairs. This further decreases training time, but the PSNR suffers due to $\mathcal{D}$ only rectifies the coarse LL subbands. Incorporating $\mathcal{U}$ to rectify HF subbands leads to the best results, at the cost of some minor additional training time.
@@ 6 Conclusion
We introduce WaveletGaussian, a sparse-view 3D Gaussian object reconstruction framework that leverages a wavelet-domain diffusion model for novel view repairs. The switch from RGB to lower-resolution wavelet domain significantly reduces overall training time, while enabling frequency-separated repairs with no information loss, as supported by experimental results.

### v3

@@ 4.2 Quantitative Results
We present the quantitative results in Table 1. Generally, compared to the closest baseline, GaussianObject [20], our proposed method achieves a 0.3-0.5 dB increase in PSNR and cuts the overall training time roughly by 40%.
@@ 4.3 Ablation Studies
Table 2 presents ablation results. Firstly, we replace the LOO strategy, utilized by the baseline, with two variations of the random masking strategy. Different from the Online RM strategy presented in Section 3.3, the Offline RM strategy does not incorporate drifting masks. The former achieves better PSNR because the more diverse corruption patterns make $\mathcal{D}$ more robust. Both strategy outperform LOO in training time due to training a single $\mathcal{G}_{d}$, and without performance reductions. Having incorporated ORM, we then use wavelet diffusion (“wavelet-$\mathcal{D}$”) for novel view repairs. This further decreases training time, but the PSNR suffers because $\mathcal{D}$ only rectifies the coarse LL subbands. Incorporating $\mathcal{U}$ to rectify HF subbands leads to the best results, at the cost of some minor additional training time.
@@ 5 Conclusion
We introduce WaveletGaussian, a sparse-view 3D Gaussian object reconstruction framework that leverages a wavelet-domain diffusion model for novel view repairs. The switch from RGB to lower-resolution wavelet domain significantly reduces overall training time, while enabling frequency-separated repairs without performance degradation, as supported by experimental results.
@@
Acknowledgements. The first author was supported by the Vingroup Science and Technology Scholarship Program for Overseas Study for Master’s and Doctoral Degrees.